OpenStudy (anonymous):

true or false for angle Θ between 0 and π/2. if tanΘ=(1/3), then sinΘ = (√10)/10.

OpenStudy (anonymous):

Theta = arctan(1/3) = 0.32175... Theta = arcsin(sqrt(10)/10) = 0.32175... Answer is true

OpenStudy (anonymous):

this one is a tricky and easy way to solve the problem

OpenStudy (anonymous):

i think i need the easy way because we haven't covered arctan yet lol

OpenStudy (anonymous):

arctan = tan^(-1) (1/3)

OpenStudy (anonymous):

or 1/tan(1/3)

OpenStudy (anonymous):

is that the only way to solve it?

OpenStudy (anonymous):

i tried to solve for cosΘ and got 3√10/10.

OpenStudy (anonymous):

well i tryed to go your way but i cant go anywhere... Im afraid the only solution i can gave you to this problem is the one i already gave you